OpenStudy (anonymous):

Find all the real square roots of 0.0036

OpenStudy (ashleyisakitty):

Do you know what the square root of 36 is?

OpenStudy (anonymous):

yeah, 6

OpenStudy (ashleyisakitty):

Awesome, all you have to do is add a point, and you got your answer! .6

OpenStudy (anonymous):

thats not an option

OpenStudy (anonymous):

A. 0.06 B. 0.6 and –0.6 C. 0.006 and –0.006 D. 0.06 and –0.06

OpenStudy (anonymous):

\[.0036=\frac{36}{10000}\] might help

OpenStudy (anonymous):

there are two square roots, the positive one, and the negative one take the square root top and bottom, two answers will be \(\frac{6}{100}\) and \(-\frac{6}{100}\)

OpenStudy (anonymous):

okay, so its B?

OpenStudy (raden):

sqrt (0.0036) = sqrt(36/10000) = sqrt(36)/sqrt(10000) = 6/100 = 0.06